Please help me to solve this linear equation: Three times a number is 3 less than twice a second number, while 7 times the first number is 5 more than 4 times the second number. find the numbers.
Answer by lwsshak3(6766)

Please help me to solve this linear equation: Three times a number is 3 less than twice a second number, while 7 times the first number is 5 more than 4 times the second number. find the numbers.
**
let x=first number
let y=second number
..
3x=2y-3
7x=4y+5
..
6x=4y-6
7x=4y+5
subtract
x=11
2y=3x+3=33+3=36
y=18
ans:
first number: 11
second number: 18